Formulate the expression to calculate normalization constant of wave function


1
Expert's answer
2021-02-07T19:10:17-0500

Because the particle must be found somewhere between x=-∞ and x=+∞ the wave function must obey the normalization condition


"1=\int ^{\infty}_{-\infty}\psi^*\psi dx"